40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jun 7 01:46:05 2012, yuta, Summary, Green Locking, Y green beat - found it!! CIMG1504.JPG
    Reply  Thu Jun 7 02:59:31 2012, yuta, Update, Green Locking, Y green beat - found it!! 
Message ID: 6777     Entry time: Thu Jun 7 02:59:31 2012     In reply to: 6775
Author: yuta 
Type: Update 
Category: Green Locking 
Subject: Y green beat - found it!! 

Summary:
  I found the big green beat note for the Y arm. The alignment of the green optics on the PSL table was crappy.

What I did:
  1. By adjusting PSL laser temperature, I found tiny beat note when

  PSL laser temperature on display: 31.35 deg C (PSL HEPA 100%)
  C1:PSL-FSS_SLOWDC = 1.75

and

  PSL laser temperature on display: 33.21 deg C (PSL HEPA 100%)
  C1:PSL-FSS_SLOWDC = -6.82

Y end laser temperature settings are fixed as follows during the measurement.

  Y end laser "T+": 34.049 deg C
  Y end laser "ADJ": 0
  Y end laser measured temperature: 34.13 deg C (*)
  C1:GCY-SLOW_SERVO2_OFFSET = 29845

Bryan's formula (swapped one; see elog #6746),  suggests the paring

  (Yend laser temp, PSL laser temp) = (34.13 deg C, 31.09 deg C).

  2. Checked that beat PD is working by swapping the beat PDs for Y arm and X arm.

  3. Checked that the mode-matching of the two beams, one from Y arm and the other from PSL, is OK by moving mode-matching lens and measuring the beam spot size at near/far field are the same.

  4. When checking the beam spot size at far field(~ 1 m from the BS), I noticed the relative beam tilt by ~ 1 mrad. We aligned them few days ago, but I think the green beam from the Y arm has shifted. Of course we align IR to the Y arm first, but we difinitely need dither servo or A2L for the arm, too.

  5. As soon as aligning the PSL green optics near the BS, I found a large beat note. The measured amplitude was ~ -26 dBm, without any amplifiers after the PD.

  Currently the measured green beam power onto the beat PD from Y end is 75 uW and from PSL is 92 uW. So the calculated beat amplitude will be ~ -10 dBm (see calculation in elog #6746). So there is about 84% loss. Anyway, I will go on to the mode scan.

ELOG V3.1.3-